# GaussianEditor: Editing 3D Gaussians Delicately with Text Instructions

### v1

@@ Abstract
Recently, impressive results have been achieved in 3D scene editing with text instructions based on a 2D diffusion model. However, current diffusion models primarily generate images by predicting noise in the latent space, and the editing is usually applied to the whole image, which makes it challenging to perform delicate, especially localized, editing for 3D scenes. Inspired by recent 3D Gaussian splatting, we propose a systematic framework, named GaussianEditor, to edit 3D scenes delicately via 3D Gaussians with text instructions. Benefiting from the explicit property of 3D Gaussians, we design a series of techniques to achieve delicate editing. Specifically, we first extract the region of interest (RoI) corresponding to the text instruction, aligning it to 3D Gaussians. The Gaussian RoI is further used to control the editing process. Our framework can achieve more delicate and precise editing of 3D scenes than previous methods while enjoying much faster training speed, i.e. within 20 minutes on a single V100 GPU, more than twice as fast as Instruct-NeRF2NeRF (45 minutes – 2 hours) [^1] The editing time varies in different scenes according to the scene structure complexity..
@@ 1 Introduction
Neural radiance field methods [29, 46, 31, 2, 3, 6, 43] have shown great power in representing 3D scenes and synthesizing novel-view images. Past years have witnessed the rapid development of NeRF and its variants, from both quality and efficiency perspectives. Editing a pre-trained NeRF model becomes a promising way to effectively edit 3D scenes. Represented by Instruct-NeRF2NeRF [11], researchers propose to use the image-conditioned 2D diffusion model, e.g. InstructPix2Pix [4], to edit 3D scenes simply with text instructions. Notable results have been achieved as real scenes can be changed following the given text instruction. However, as current 2D diffusion models have the limited ability to localize editing regions, it is hard to generate delicate edited scenes as the unwanted regions are usually changed with diffusion models. Even though some works [30] propose to constrain the editing region on the edited 2D images, the editing region is still hard to accurately localize and apply to the 3D representation. Besides, NeRF-based methods [49, 9] bear coupling effects between different spatial positions, e.g. different points are queried from the same MLP field (for implicit representations) or voxel vertices (for explicit representations).
Recent 3D Gaussian Splatting [18] (3D-GS) has been a groundbreaking work in the field of radiance field, which is the first to achieve a real sense of real-time rendering while enjoying high rendering quality and training speed. Besides its efficiency, we further notice its natural explicit property. 3D-GS has a great advantage for editing tasks as each 3D Gaussian exists individually. It is easy to edit 3D scenes by directly manipulating 3D Gaussians with desired constraints applied.
Aiming at editing 3D scenes delicately, we propose to represent the scene with 3D Gaussians, which can be edited with text instructions, and name our method as GaussianEditor. To achieve precise control for editing regions, we divide our method into three main parts. The first one is the region of interest (RoI) extraction from the given text instruction. The instruction may be complex or indirect while this module helps extract the key words matching the RoI for editing. The second part aligns the extracted text RoI to the 3D Gaussian space through the image space, where a grounding segmentation model can be applied. The last part is to edit the original 3D Gaussians delicately with constraints in the obtained 3D Gaussian RoI. With the above processes, the region for editing can be precisely localized simply from text instructions, which are used for constraining the 3D Gaussian updating to obtain a delicately edited new 3D scene. Besides, we enable interfaces for users to introduce more exact instructions for more delicate editing, e.g. Gaussian point selecting and 3D boxes for modifying the editing regions [^2] These additional instructions are applied to generate the man with two different edited half faces in Fig. 1..
@@ 2 Related Work
2D Image Editing with Diffusion Models
@@
By virtue of advancements in diffusion model technology [13, 44], numerous generative models [41] have attained remarkably impressive outcomes in the realm of image synthesis. Recent progressions in diffusion models have showcased their adeptness in fabricating lifelike images from arbitrary textual inputs [8, 14, 40, 42, 45]. Harnessing the robust semantic comprehension and image generation capabilities of foundational diffusion models, an escalating number of research explorations are currently employing diffusion models as a fundamental framework for implementing text-based image editing functionalities [33, 37, 38, 41]. Some of these methodologies necessitate the manual provision of captions for both the original and edited images [12], while others mandate specific scenario-based training for optimization [39]. These requisites have rendered it arduous for ordinary users to avail themselves of such techniques. Expanding upon this foundation, iP2P [4] introduces instruction-based capabilities to image editing, enabling users to simply input an image and apprise the model of the desired alterations. This user-friendly approach facilitates the democratization of image editing in a more accessible manner.
@@
Figure 2: Our framework, named GaussianEditor, consists of three key steps. First, the scene generation ${\mathcal{F}}$ was performed to get the scene description ${\mathcal{T}}_{dsc}$ from original 3D Gaussians ${\mathcal{G}}_{ori}$. The description ${\mathcal{T}}_{dsc}$ and the text instruction ${\mathcal{T}}$ provided by the user are then combined using a template ${\mathcal{T}}_{tmp}$ to get user message ${\mathcal{T}}_{usr}$. The ${\mathcal{T}}_{usr}$ and a pre-defined prompt ${\mathcal{P}}_{e}$ are fed into an LLM assistant ${\mathcal{A}}$ to obtain the instruction RoI ${\mathcal{T}}_{RoI}$. Second, a grounding segmentation model ${\mathcal{S}}$ is used to convert ${\mathcal{T}}_{RoI}$ to image RoI ${\mathcal{I}}_{RoI}$, which is then lifted to 3D Gaussians RoI ${\mathcal{G}}_{RoI}$ by RoI lifting ${\mathcal{H}}$, where additional user instructions ${\mathcal{O}}$ can be incorporated. Third, following the user instruction ${\mathcal{T}}$, rendered image ${\mathcal{I}}_{rd}$ from randomly chosen views is edited by a 2D diffusion model ${\mathcal{D}}$. The loss between ${\mathcal{I}}_{rd}$ and edited one ${\mathcal{I}}_{ed}$ is calculated. Finally, gradient backpropagation and optimization are performed within the Gaussian RoI ${\mathcal{G}}_{RoI}$ to get the edited scene ${\mathcal{G}}_{ed}$.
3D Scene Editing of Radiance Fields
@@
3D Scene Editing of Radiance Fields has gained significant popularity as a recent research direction [25, 28, 15, 48, 20, 47, 1, 10, 34, 24, 22, 53, 54, 52, 23]. The main objective of such methods is to enable the manipulation of both the geometry and appearance of 3D scene representations. However, editing such scenes poses inherent challenges due to the implicit nature of traditional NeRF representations, which lack precise localization capabilities. As a result, previous works have primarily focused on achieving global style transformations of 3D scenes [49, 7, 16, 17, 32, 56, 51]. While some efforts have been made towards object-centric scene editing[57], keeping the background unchanged has been a persistent challenge. For example, a recently proposed work, Instruct-NeRF2NeRF [11], implements text instruction-controlled 3D scene editing capabilities, achieving excellent editing effects while maintaining user-friendliness. However, it relies on the editing effect of 2D images, which may cause global changes to the 3D scene. A subsequent work [30] attempts to compute the relevance map between edited and unedited images to localize the editing area. The relevance map may be unreliable when the 2D IP2P [4] model fails. Other efforts [23] attempt to rely on the 3D coordinates entered by the user to decide the editing area. The introduction of 3D Gaussians [18] has provided an opportunity to address this limitation. Its explicit 3D representation methodology enables accurate selection and manipulation of editing areas. By incorporating LLMs, the whole process can be more automated.
@@ 3 Method
In this section, we first review 3D representation methods in Sec. 3.1. Subsequently, in Sec. 3.2, we overview our proposed approach, which mainly includes three modules. Sec. 3.3 delves into the precise Region of Interest (RoI) extraction of text instructions, while Sec. 3.4 introduces how to align the instruction RoI with 3D Gaussians. Finally, Sec. 3.5 describes the delicate editing process within the obtained Gaussian RoI.
@@ 3.1 Preliminaries
$C=\sum_{i\in{\mathcal{N}}}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $\sigma_{i}=\alpha_{i}e^{-\frac{1}{2}(x_{i})^{T}\Sigma^{-1}(x_{i})}$ represents the influence of the Gaussian to the image pixel and $x_{i}$ is the distance between the pixel and the center of the i-th Gaussian.
@@ 3.2 Overall Framework
Given a group of 3D Gaussians ${\mathcal{G}}_{ori}$ for a scene and a text instruction ${\mathcal{T}}$ for editing, our Gaussian editor ${\mathcal{E}}$ can edit the 3D Gaussians delicately into a new one, denoted as ${\mathcal{G}}_{ed}$, with the guidance of the instruction. The whole process can be formulated as
${\mathcal{G}}_{ed}={\mathcal{E}}({\mathcal{G}}_{ori},{\mathcal{T}}).$ (2)
Fig. 2 illustrates the overall framework of our approach, which consists of three main steps. First, the region of interest (RoI) is extracted from the text instruction. In this step, we employ a multimodal model ${\mathcal{C}}$ to generate the scene description ${\mathcal{T}}_{dsc}$ from the rendered image ${\mathcal{I}}_{rd}$ of original 3D Gaussians $\mathcal{G}_{ori}$. We then input the scene description ${\mathcal{T}}_{dsc}$ and text instruction ${\mathcal{T}}$ into a large language model assistant ${\mathcal{A}}$ to determine where we should make edits in the scene. The output of this step is referred to as the instruction Region of Interest (RoI) $\mathcal{T}_{RoI}$.
The next step is the 3D Gaussian RoI alignment. We use a grounding segmentation model ${\mathcal{S}}$ to convert the RoI from text space, i.e. ${\mathcal{T}}_{RoI}$, to the image space, i.e. ${\mathcal{I}}_{RoI}$. Then the image RoI ${\mathcal{I}}_{RoI}$ is lifted to the RoI of 3D Gaussians ${\mathcal{G}}_{RoI}$ through training. The Gaussian RoI allows us to precisely control the regions where edits will be applied.
The last step is delicate editing within the Gaussian RoI. In this step, we randomly sample the view to obtain the rendered image ${\mathcal{I}}_{rd}$. A 2D diffusion model ${\mathcal{D}}$ is used to perform the editing process on the rendered image ${\mathcal{I}}_{rd}$, with the user instruction ${\mathcal{T}}$ and the original image ${\mathcal{I}}_{ori}$ as conditions. The resulting edited image is denoted as ${\mathcal{I}}_{ed}$. Subsequently, we calculate the loss between ${\mathcal{I}}_{ed}$ and ${\mathcal{I}}_{rd}$ and make gradient back-propagation within ${\mathcal{G}}_{RoI}$. This implies that only the regions specified by the RoI can receive corresponding gradients during the back-propagation process. Finally, optimization is executed based on these gradients. Through several rounds of iterative optimization, the final optimized scene representation ${\mathcal{G}}_{ed}$ is obtained.
@@ 3.3 RoI Extraction of Text Instruction
The instruction RoI is extracted for the editing regions from both the 3D scene representation ${\mathcal{G}}_{ori}$ and the text instruction ${\mathcal{T}}$ provided by the user. To achieve this, we employ a multimodal model ${\mathcal{C}}$ in conjunction with the large language model assistant ${\mathcal{A}}$. The first step is the scene description generation ${\mathcal{F}}$, which aims to get the text description ${\mathcal{T}}_{dsc}$ from 3D Gaussians ${\mathcal{G}}_{ori}$:
${\mathcal{T}}_{dsc}={\mathcal{F}}({\mathcal{G}}_{ori}).$ (3)
By leveraging the technique of differentiable splatting as shown in Eq. 1, a set of 2D image samples ${\mathcal{I}}_{rd}$ are generated, the quantity of which is recorded as ${\mathcal{M}}$. These sampled images ${\mathcal{I}}{sp}=\{{\mathcal{I}}_{1},{\mathcal{I}}_{2},...,{\mathcal{I}}_{\mathcal{M}}\}$, are then inputted into a multimodal model ${\mathcal{C}}$ to generate corresponding text descriptions ${\mathcal{T}}_{img}=\{{\mathcal{T}}_{1},{\mathcal{T}}_{2},...,{\mathcal{T}}_{\mathcal{M}}\}$:
${\mathcal{T}}_{k}={\mathcal{C}}({\mathcal{P}}_{c},{\mathcal{I}}_{k}),k\in~{}{\mathcal{M}},$ (4)
where ${\mathcal{P}}_{c}$ is a prompt given by the user, such as “What is the content of the image”, to multimodal model ${\mathcal{C}}$ for getting precise description. Subsequently, these descriptions are fed into a large language model, which works as a scene understanding assistant ${\mathcal{A}}$. The assistant ${\mathcal{A}}$ is specifically instructed by a prompt ${\mathcal{P}}_{m}$, to merge descriptions of diverse views into one detailed scene description ${\mathcal{T}}_{dsc}$:
${\mathcal{T}}_{dsc}={\mathcal{A}}({\mathcal{P}}_{m},{\mathcal{T}}_{img}).$ (5)
The process of the scene description generation ${\mathcal{F}}$ is shown in Fig. 3.
Next, large language model assistant ${\mathcal{A}}$ is used to extract the instruction RoI. Initially, we set a prompt ${\mathcal{P}}_{e}$ to instruct ${\mathcal{A}}$ about the task for assisting in instruction RoI extraction. Then the scene description ${\mathcal{T}}_{dsc}$ and the user instruction ${\mathcal{T}}$ are combined with a predefined template ${\mathcal{T}}_{tmp}$: “Text description: ${\mathcal{T}}_{dsc}$ Edit Instruction: ${\mathcal{T}}$ Answer:” to form the user message ${\mathcal{T}}_{usr}$=${\mathcal{T}}_{tmp}({\mathcal{T}}_{dsc},{\mathcal{T}})$. During the actual execution, we input both the prompt ${\mathcal{P}}_{e}$ and the user message ${\mathcal{T}}_{usr}$ into ${\mathcal{A}}$ to obtain the instruction RoI ${\mathcal{T}}_{RoI}$:
${\mathcal{T}}_{RoI}={\mathcal{A}}({\mathcal{P}}_{e},{\mathcal{T}}_{usr}).$ (6)
@@
To confine the 3D editing region within the instruction RoI, 3D Gaussian RoI ${\mathcal{G}}_{RoI}$ is aligned with the text RoI ${\mathcal{T}}_{RoI}$ in this step. First, The RoI in the text space is transformed into the image space via a grounding segmentation model ${\mathcal{S}}$:
${\mathcal{I}}_{RoI}={\mathcal{S}}({\mathcal{I}}_{rd},{\mathcal{T}}_{RoI}),$ (7)
where ${\mathcal{I}}_{rd}$ is rendered image of the original scene ${\mathcal{G}}_{ori}$.
Then we lift the the RoI ${\mathcal{I}}_{RoI}$ in the image space to 3D Gaussian ${\mathcal{G}}_{RoI}$ through training. To achieve this, an additional RoI attribute $r\in{\mathbb{R}}^{1}$ was added to 3D Gaussian $g_{i}=\{\mu_{i},\Sigma_{i},c_{i},\alpha_{i},r_{i}\}$. $r$ is initialized to 0, which means it is not in the Gaussians RoI, and 1 means it is inside the RoI. The set of $r$ is denoted as ${\mathcal{R}}\in{\mathbb{R}}^{{\mathcal{N}},1}$, where the ${\mathcal{N}}$ is the number of 3D Gaussians ${\mathcal{G}}_{ori}$.
Then the color $c_{i}$ in Eq. 1 was rewritten with $r_{i}$ to get the rendered RoI ${\mathcal{I}}_{RoI}^{rd}$:
${\mathcal{I}}_{RoI}^{rd}=\sum_{i\in{\mathcal{N}}}r_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (8)
Taking inspiration from SA3D [5], to get the trained Gaussians RoI ${\mathcal{G}}_{RoI}^{tr}$, we adopt a similar loss function to supervise the training process:
$\mathcal{L}_{proj}=\lambda_{1}\sum({\mathcal{I}}_{RoI}^{rd}\cdot{\mathcal{I}}_{RoI})+\lambda_{2}\sum((1-{\mathcal{I}}_{RoI}^{rd})\cdot{\mathcal{I}}_{RoI}),$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. Additionally, the user can modify the trained Gaussian RoI ${\mathcal{G}}_{RoI}^{tr}$ by giving added Gaussian RoI ${\mathcal{G}}_{RoI}^{add}$, deleted Gaussian RoI ${\mathcal{G}}_{RoI}^{del}$ and 3D box ${\mathcal{B}}_{3D}$:
${\mathcal{G}}_{RoI}=({\mathcal{G}}_{RoI}^{tr}\cup~{}{\mathcal{G}}_{RoI}^{add}-{\mathcal{G}}_{RoI}^{del})\cap~{}{\mathcal{B}}_{3D},$ (10)
${\mathcal{G}}_{RoI}^{add}$ represents the 3D Gaussians user wants to edit, ${\mathcal{G}}_{RoI}^{del}$ means 3D Gaussians user wants to keep from editing, ${\mathcal{B}}_{3D}$ is the coordinates of 3D cuboid it limits RoI to inside the box. ${\mathcal{G}}_{RoI}$ is the aligned RoI with the text RoI. For example, when editing the left face of the man in Fig. 1, grounding segmentation failed to ground “left face”, instead, it grounded the whole face. In this scenario, the user can use the interactive interface to set the right face as ${\mathcal{G}}_{RoI}^{del}$ or enter the rectangular box where the left face is located as ${\mathcal{B}}_{3D}$. The lifting process ${\mathcal{H}}$ can be represented as:
${\mathcal{G}}_{RoI}={\mathcal{H}}({\mathcal{I}}_{RoI},{\mathcal{O}}),$ (11)
where ${\mathcal{O}}=\{{\mathcal{G}}_{RoI}^{add},{\mathcal{G}}_{RoI}^{del},{\mathcal{B}}_{3D}\}$ is additional user instructions.

### v2

@@ Abstract
Recently, impressive results have been achieved in 3D scene editing with text instructions based on a 2D diffusion model. However, current diffusion models primarily generate images by predicting noise in the latent space, and the editing is usually applied to the whole image, which makes it challenging to perform delicate, especially localized, editing for 3D scenes. Inspired by recent 3D Gaussian splatting, we propose a systematic framework, named GaussianEditor, to edit 3D scenes delicately via 3D Gaussians with text instructions. Benefiting from the explicit property of 3D Gaussians, we design a series of techniques to achieve delicate editing. Specifically, we first extract the region of interest (RoI) corresponding to the text instruction, aligning it to 3D Gaussians. The Gaussian RoI is further used to control the editing process. Our framework can achieve more delicate and precise editing of 3D scenes than previous methods while enjoying much faster training speed, i.e. within 20 minutes on a single V100 GPU, more than twice as fast as Instruct-NeRF2NeRF (45 minutes – 2 hours) [^1] The editing time varies in different scenes according to the scene structure complexity.. The project page is at https://GaussianEditor.github.io.
@@ 1 Introduction
Neural radiance field methods [29, 46, 31, 2, 3, 6, 43] have shown great power in representing 3D scenes and synthesizing novel-view images. Past years have witnessed the rapid development of NeRF and its variants, from both quality and efficiency perspectives. Editing a pre-trained NeRF model becomes a promising way to edit 3D scenes. Represented by Instruct-NeRF2NeRF [11], researchers propose to use the image-conditioned 2D diffusion model, e.g. InstructPix2Pix [4], to edit 3D scenes simply with text instructions. Notable results have been achieved as real scenes can be changed following the text instruction. However, current 2D diffusion models face challenges in accurately localizing editing regions, which hinders the generation of finely edited scenes due to the change of unintended regions. Even though some works [30] propose to constrain the editing region on edited 2D images, the editing region is not accurately localized and hard to apply to the 3D representation. Besides, NeRF-based methods [49, 9] bear coupling effects between different spatial positions, e.g. different points are queried from the same MLP field (for implicit representations) or voxel vertices (for explicit representations).
Recent 3D Gaussian Splatting [18] (3D-GS) has been a groundbreaking work in the radiance field, which is the first to achieve a real sense of real-time rendering while enjoying high rendering quality and training speed. Besides its efficiency, we further notice its natural explicit property. 3D-GS has a great advantage for editing tasks as each 3D Gaussian exists individually. Editing 3D scenes by directly manipulating 3D Gaussians with desired constraints is easy.
Aiming at editing 3D scenes delicately, we propose to represent the scene with 3D Gaussians, which can be edited with text instructions, and name our method as GaussianEditor. GaussianEditor is divided into three main parts to achieve precise control for editing regions. The first is the region of interest (RoI) extraction from the given text instruction. The instruction may be complex or indirect while this module helps extract the keywords matching the RoI for editing. The second part aligns the extracted text RoI to the 3D Gaussian space through the image space, where a grounding segmentation module is applied. The last part is to edit the original 3D Gaussians delicately with constraints in the obtained 3D Gaussian RoI. With the above processes, the region for editing can be precisely localized simply from text instructions, which constrains the 3D Gaussian updating to obtain a delicately edited new 3D scene. Besides, we enable interfaces for users to introduce more exact instructions for more delicate editing, e.g. Gaussian point selecting and 3D boxes for modifying the editing regions [^2] These additional instructions are applied to generate the man with two different edited half faces in Fig. 1..
@@ 2 Related Work
2D Image Editing with Diffusion Models.
@@
Advancements in diffusion model technology [13, 44], have led to numerous generative models [41] achieving impressive outcomes in image synthesis. Recent developments in diffusion models have demonstrated their ability to create lifelike images from arbitrary textual inputs [8, 14, 40, 42, 45]. Harnessing the robust semantic comprehension and image generation capabilities of foundational diffusion models, an escalating number of research explorations are currently employing diffusion models as a fundamental framework for implementing text-based image editing functionalities [33, 37, 38, 41]. Some of these methodologies necessitate the manual provision of captions for both the original and edited images [12], while others mandate specific scenario-based training for optimization [39]. These requisites have rendered it arduous for ordinary users to avail themselves of such techniques. Expanding upon this foundation, iP2P [4] introduces instruction-based capabilities to image editing, enabling users to simply input an image and apprise the model of the desired alterations. This user-friendly approach facilitates the democratization of image editing in a more accessible manner.
@@
Figure 2: Our framework, named GaussianEditor, consists of three key steps. First, a module ${\mathcal{M}}_{Desc}$ is used to get the description of the input scene, which is put to an LLM assistant ${\mathcal{M}}_{LLM}$ with the text instruction ${\mathcal{T}}$ provided by the user to obtain the text RoI ${\mathcal{T}}_{RoI}$. Second, a grounding segmentation module ${\mathcal{M}}_{Seg}$ is used to convert ${\mathcal{T}}_{RoI}$ to image RoI ${\mathcal{I}}_{RoI}$, which is then lifted to 3D Gaussians RoI ${\mathcal{G}}_{RoI}$ by RoI lifting ${\mathcal{M}}_{Lift}$, where additional user instructions ${\mathcal{O}}$ can be incorporated. Third, following the user instruction ${\mathcal{T}}$, rendered image ${\mathcal{I}}_{render}$ from randomly chosen views is edited by a diffusion model ${\mathcal{M}}_{DM}$. The loss between ${\mathcal{I}}_{render}$ and edited one ${\mathcal{I}}_{edit}$ is calculated. Finally, gradient backpropagation and optimization are performed within the Gaussian RoI ${\mathcal{G}}_{RoI}$ to get the edited scene ${\mathcal{G}}_{edit}$.
3D Scene Editing of Radiance Fields.
@@
3D Scene Editing of Radiance Fields has become a popular research direction [25, 28, 15, 48, 20, 47, 1, 10, 34, 24, 22, 53, 54, 52, 23]. These methods aim to manipulate the geometry and appearance of 3D scene representations. However, editing such scenes poses challenges due to the implicit nature of traditional NeRF representations, which lack precise localization capabilities. As a result, previous works have primarily focused on achieving global style transformations of 3D scenes [49, 7, 16, 17, 32, 58, 51]. While some efforts have been made towards object-centric scene editing[59], keeping the background unchanged has been a persistent challenge. For example, the recently proposed Instruct-NeRF2NeRF [11] implements text instruction-controlled 3D scene editing, achieving excellent editing effects while maintaining user-friendliness. However, it relies on the editing effect of 2D images, which may cause global changes to the 3D scene. A subsequent work [30] attempts to compute the relevance map between edited and unedited images to localize the editing area. The relevance map may be unreliable when the 2D IP2P [4] model fails. Other efforts [23] rely on the user-entered 3D coordinates to determine the editing area. The introduction of 3D Gaussians [18] has provided an opportunity to address this limitation. Its explicit 3D representation enables accurate selection and manipulation of editing areas. By incorporating LLMs, the whole process can be more automated.
@@ 3 Method
In this section, we first review 3D representation methods in Sec. 3.1. Subsequently, in Sec. 3.2, we overview our proposed approach, which mainly includes three modules. Sec. 3.3 delves into the precise Region of Interest (RoI) extraction of text instructions, using scene description generation module ${\mathcal{M}}_{Desc}$ and LLM assistant ${\mathcal{M}}_{LLM}$. Sec. 3.4 introduces how to align the instruction RoI with 3D Gaussians, using grounding segmentation module ${\mathcal{M}}_{Seg}$ and RoI lifting module ${\mathcal{M}}_{Lift}$. Finally, Sec. 3.5 describes the delicate editing process within the obtained Gaussian RoI, using text to image diffusion model ${\mathcal{M}}_{DM}$.
@@ 3.1 Preliminaries
$C=\sum_{i\in{\mathcal{N}}}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j}),$ (1)
where $\sigma_{i}=\alpha_{i}e^{-\frac{1}{2}(x_{i})^{T}\Sigma^{-1}(x_{i})}$ represents the influence of the Gaussian to the image pixel and $x_{i}$ is the distance between the 3D point and the center of the i-th Gaussian.
@@ 3.2 Overall Framework
Given a group of 3D Gaussians ${\mathcal{G}}_{input}$ for an input scene and a text instruction ${\mathcal{T}}$ for editing, our Gaussian editor ${\mathcal{E}}$ can edit the 3D Gaussians delicately into a new one, denoted as ${\mathcal{G}}_{edit}$, with the guidance of the instruction. The whole process can be formulated as
${\mathcal{G}}_{edit}={\mathcal{E}}({\mathcal{G}}_{input},{\mathcal{T}}).$ (2)
Fig. 2 illustrates the overall framework of our approach, which consists of three main steps. First, the Region of Interest (RoI) is extracted from the text instruction. In this step, we employ a module named scene description generation ${\mathcal{M}}_{Dexcription}$ to get the description of the input scene. We then input the scene description ${\mathcal{T}}_{scene}$ and text instruction ${\mathcal{T}}$ into a large language model assistant ${\mathcal{M}}_{LLM}$ to determine where we should make edits in the scene. The output of this step is referred to as the instruction RoI ${\mathcal{T}}_{RoI}$.
The next step is the 3D Gaussian RoI alignment. We use a grounding segmentation module ${\mathcal{M}}_{Seg}$ to convert the RoI from text space, i.e. ${\mathcal{T}}_{RoI}$, to the image space, i.e. ${\mathcal{I}}_{RoI}$. Then the image RoI ${\mathcal{I}}_{RoI}$ is lifted to the RoI of 3D Gaussians ${\mathcal{G}}_{RoI}$ through RoI lifting module ${\mathcal{M}}_{Lift}$. The Gaussian RoI allows us to control the regions where edits will be applied precisely.
The last step is delicate editing within the Gaussian RoI. In this step, we randomly sample the view to obtain the rendered image ${\mathcal{I}}_{render}$. A 2D diffusion model ${\mathcal{M}}_{DM}$ is used to perform the editing process on the rendered image ${\mathcal{I}}_{render}$, with the user instruction ${\mathcal{T}}$ and the image ${\mathcal{I}}_{input}$ of input scene as conditions. The resulting edited image is denoted as ${\mathcal{I}}_{edit}$. Subsequently, we calculate the loss between ${\mathcal{I}}_{edit}$ and ${\mathcal{I}}_{render}$ and make gradient back-propagation within ${\mathcal{G}}_{RoI}$. This implies that only the regions specified by the RoI can receive corresponding gradients during the back-propagation process. Finally, optimization is executed based on these gradients. The final optimized scene representation ${\mathcal{G}}_{edit}$ is obtained through several rounds of iterative optimization.
@@ 3.3 RoI Extraction of Text Instruction
The instruction RoI is extracted for the editing regions from both the input 3D scene ${\mathcal{G}}_{input}$ and the text instruction ${\mathcal{T}}$ provided by the user. To achieve this, we employ a multimodal model ${\mathcal{M}}_{MM}$ in conjunction with the large language model assistant ${\mathcal{M}}_{LLM}$. The first step is the scene description generation ${\mathcal{M}}_{Desc}$, which aims to get the scene description ${\mathcal{T}}_{scene}$ from 3D Gaussians ${\mathcal{G}}_{input}$:
${\mathcal{T}}_{scene}={\mathcal{M}}_{Desc}({\mathcal{G}}_{input}).$ (3)
The process of the scene description generation ${\mathcal{M}}_{Desc}$ is shown in Fig. 3. By leveraging the technique of differentiable splatting as shown in Eq. 1, a set of 2D image samples $\{{\mathcal{I}}_{sample}\}$ are generated and then inputted into a multimodal model ${\mathcal{M}}_{MM}$ to generate corresponding text descriptions $\{{\mathcal{T}}_{sample}\}$:
${\mathcal{T}}_{sample}={\mathcal{M}}_{MM}({\mathcal{P}}_{MM},{\mathcal{I}}_{sample}),$ (4)
where ${\mathcal{P}}_{MM}$ is a prompt, such as “What is the content of the image”, for multimodal model ${\mathcal{M}}_{MM}$ to get precise description. Subsequently, these descriptions $\{{\mathcal{T}}_{sample}\}$ are fed into a large language model ${\mathcal{M}}_{LLM}$, which is specifically instructed by a prompt ${\mathcal{P}}_{merge}$ to merge descriptions of diverse views into one detailed scene description ${\mathcal{T}}_{scene}$:
${\mathcal{T}}_{scene}={\mathcal{M}}_{LLM}({\mathcal{P}}_{merge},\{{\mathcal{T}}_{sample}\}).$ (5)
After that, the scene description ${\mathcal{T}}_{scene}$ and the user instruction ${\mathcal{T}}$ are combined with a predefined template ${\mathcal{T}}_{template}$: “Text description: ${\mathcal{T}}_{scene}$ Edit Instruction: ${\mathcal{T}}$ Answer:” to form the user message ${\mathcal{T}}_{user}$=${\mathcal{T}}_{template}({\mathcal{T}}_{scene},{\mathcal{T}})$. The LLM model ${\mathcal{M}}_{LLM}$ is used to extract the instruction RoI ${\mathcal{T}}_{RoI}$ from user message ${\mathcal{T}}_{user}$ with a new prompt ${\mathcal{P}}_{extract}$:
${\mathcal{T}}_{RoI}={\mathcal{M}}_{LLM}({\mathcal{P}}_{extract},{\mathcal{T}}_{user}).$ (6)
@@
To confine the 3D editing region within the instruction RoI, 3D Gaussian RoI ${\mathcal{G}}_{RoI}$ is aligned with the text RoI ${\mathcal{T}}_{RoI}$. First, The RoI in the text space is transformed into the image space via a grounding segmentation module ${\mathcal{M}}_{Seg}$:
${\mathcal{I}}_{RoI}={\mathcal{M}}_{Seg}({\mathcal{I}}_{input},{\mathcal{T}}_{RoI}),$ (7)
where ${\mathcal{I}}_{input}$ is rendered image of the input scene ${\mathcal{G}}_{input}$.
Then we lift the the RoI ${\mathcal{I}}_{RoI}$ in the image space to 3D Gaussian ${\mathcal{G}}_{RoI}$ through training. To achieve this, an additional RoI attribute $r\in{\mathbb{R}}^{1}$ was added to 3D Gaussian $g_{i}=\{\mu_{i},\Sigma_{i},c_{i},\alpha_{i},r_{i}\}$. $r$ is initialized to 0, which means it is not in the Gaussians RoI, and 1 means it is inside the RoI. The set of $r$ is denoted as ${\mathcal{R}}\in{\mathbb{R}}^{{\mathcal{N}},1}$, where the ${\mathcal{N}}$ is the number of 3D Gaussians ${\mathcal{G}}_{input}$.
Then the color $c_{i}$ in Eq. 1 was rewritten with $r_{i}$ to get the rendered RoI ${\mathcal{I}}_{RoI}^{render}$:
${\mathcal{I}}_{RoI}^{render}=\sum_{i\in{\mathcal{N}}}r_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j}).$ (8)
Taking inspiration from SA3D [5], to get the trained Gaussians RoI ${\mathcal{G}}_{RoI}^{train}$, we adopt a similar loss function to supervise the training process:
$\mathcal{L}_{proj}=\lambda_{1}\sum({\mathcal{I}}_{RoI}^{render}\cdot{\mathcal{I}}_{RoI})+\lambda_{2}\sum((1-{\mathcal{I}}_{RoI})\cdot{\mathcal{I}}_{RoI}^{render}),$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. The $r$ in Eq. 8 is updated via $r\leftarrow r-\eta\frac{\partial\mathcal{L}_{proj}}{\partial r}$ with gradient descent, where $\eta$ denotes the learning rate. Eq. 9 encourages rendered RoI to cover the Image RoI and not exceed it. Additionally, the user can modify the trained Gaussian RoI ${\mathcal{G}}_{RoI}^{train}$ by giving added Gaussian RoI ${\mathcal{G}}_{RoI}^{add}$, deleted Gaussian RoI ${\mathcal{G}}_{RoI}^{del}$ and 3D box ${\mathcal{B}}_{3D}$:
${\mathcal{G}}_{RoI}=({\mathcal{G}}_{RoI}^{train}\cup~{}{\mathcal{G}}_{RoI}^{add}-{\mathcal{G}}_{RoI}^{del})\cap~{}{\mathcal{B}}_{3D},$ (10)
${\mathcal{G}}_{RoI}^{add}$ represents the 3D Gaussians user wants to edit, ${\mathcal{G}}_{RoI}^{del}$ means 3D Gaussians user wants to keep from editing, ${\mathcal{B}}_{3D}$ is the coordinates of 3D cuboid it limits RoI to inside the box. ${\mathcal{G}}_{RoI}$ is the aligned RoI with the text RoI. For example, when editing the left face of the man in Fig. 1, grounding segmentation failed to ground “left face”, instead, it grounded the whole face. In this scenario, the user can use the interactive interface to set the right face as ${\mathcal{G}}_{RoI}^{del}$ or enter the rectangular box where the left face is located as ${\mathcal{B}}_{3D}$. The lifting process ${\mathcal{M}}_{Lift}$ can be represented as:
${\mathcal{G}}_{RoI}={\mathcal{M}}_{Lift}({\mathcal{I}}_{RoI},{\mathcal{O}}),$ (11)
where ${\mathcal{O}}=\{{\mathcal{G}}_{RoI}^{add},{\mathcal{G}}_{RoI}^{del},{\mathcal{B}}_{3D}\}$ is optional instructions.